OpenStudy (jessicaprieto):

If it takes 4/7 of a ton of concrete to complete 7/8 of bridge, how much is needed to complete the entire bridge

OpenStudy (joshtray):

If 4/7 of a ton is needed to complete 7/8 of a bridge, 1/8 of a bridge will need 4/7 X 1/7 which is 4/49 tons of concrete will be used. So to complete 8/8 which is the whole bridge, you will need 4/49 X 8 which is 32/49 of a ton of concrete

OpenStudy (joshtray):

If you dont understand I could explain more

OpenStudy (jessicaprieto):

If you can break it down for me that would be great.

OpenStudy (calculusxy):

\[\large \frac{\frac{4}{7} \text{ }\text{ton}}{\frac{7}{8} \text{ } \text{of a bridge}} = \frac{x}{1}\]

OpenStudy (calculusxy):

Making a proportion and trying to solve for x would be useful.

OpenStudy (mathstudent55):

You can also use a proportion. x is the amount needed for the entire bridge. 4/7 ton is to 7/8 bridge as x is to 1 bridge \(\large \dfrac{\frac{4}{7}}{\frac{7}{8}} = \dfrac{x}{1} \) Cross multiply: \(\dfrac{7}{8}x = \dfrac{4}{7} \times 1\) \(\dfrac{7}{8}x = \dfrac{4}{7} \) You have \(\dfrac{7}{8}x\), and you want x, so multiply both sides by the reciprocal of the fraction \(\dfrac{8}{7} \times \dfrac{7}{8}x = \dfrac{8}{7} \times \dfrac{4}{7} \) \(x = \dfrac{32}{49} \) Answer: \(\dfrac{32}{49}\) of a ton is needed to complete the entire bridge

OpenStudy (joshtray):

Sorry I was quite busy

OpenStudy (joshtray):

But their explanation is correct

OpenStudy (jessicaprieto):

No problem - This makes it very simple. Thank you very much